### v1

# AiXpand AI OS - Decentralized ubiquitous computing MLOps execution engine

Abstract
Over the past few years, ubiquitous, or pervasive computing has gained popularity as the primary approach for a wide range of applications, including enterprise-grade systems, consumer applications, and gaming systems. Ubiquitous computing refers to the integration of computing technologies into everyday objects and environments, creating a network of interconnected devices that can communicate with each other and with humans. By using ubiquitous computing technologies, communities can become more connected and efficient, with members able to communicate and collaborate more easily. This enabled interconnectedness and collaboration can lead to a more successful and sustainable community. The spread of ubiquitous computing, however, has emphasized the importance of automated learning and smart applications in general. Even though there have been significant strides in Artificial Intelligence and Deep Learning, large scale adoption has been hesitant due to mounting pressure on expensive and highly complex cloud numerical-compute infrastructures. Adopting, and even developing, practical machine learning systems can come with prohibitive costs, not only in terms of complex infrastructures but also of solid expertise in Data Science and Machine Learning. In this paper we present an innovative approach for low-code development and deployment of end-to-end AI cooperative application pipelines. We address infrastructure allocation, costs, and secure job distribution in a fully decentralized global cooperative community based on tokenized economics.
1 Introduction
1.1 From Web 1.0 to 2.0 and finally to 3.0
Pervasive computing applications have been a part of our lives in various forms for quite some time - from smartwatches to smartphones, from smart homes to smart traffic lights. Although some argue that ubiquitous or pervasive computing is equivalent to decentralization [1] [2] [3], most systems still depend on intricate proprietary cloud systems. The concept of fog computing [4] involves the cloud infrastructure handling the heavy lifting of hardware and logic, while edge devices manage IoT data acquisition, thin client user interaction, and basic edge processing.
Besides this, if we examine the Internet in terms of evolution, we can suggest that as the transition from Web 1.0, which was read-only, to Web 2.0 introduced a massive content explosion via creators, influencers, and advocates, the logical and natural progression would be for Web 3.0 to enable transparency and democratization. The current practice and theory have shown that decentralization [5] [6], can bring significant benefits to human society. Blockchain-based technologies [7] [8] [9] have the potential to democratize data, enable peer-to-peer compute sharing, provide affordable go-to-market options for content creators and app developers. These technologies offer secure, transparent, trustless, and low-cost micro-transactions.
Artificial Intelligence (AI), machine (deep) learning in particular, is the most important catalyst for the development of intelligent applications and pervasive computing. In the era of ChatGPT, people acknowledge the transformative power of Artificial Intelligence, but its adoption is slow due to the high costs associated with developing and running AI-enabled applications.
The primary focus in the last decade has been on developing and deploying cloud-based AI applications, largely due to the efforts of major players, such as FAANG [10], to somehow make high-performance GPUs accessible to the broader audience. Despite the efforts made, there are several problems that remain unresolved or only partially resolved, which hinder the widespread adoption of AI applications and systems. These issues arise from the high cost of AI expertise and the extensive range of skills required for developing and deploying production-grade AI systems. Additionally, the significant expenses and energy consumption of GPU computing infrastructure contribute to substantial carbon footprints. To overcome these challenges, designing and implementing deep neural models for practical use cases requires a combination of scientific and business expertise, along with practical experience.
All in all, it appears that the creation of AI applications is still far from being a truly democratized process, and several problems must be addressed before it becomes a horizontal approach that is widely accessible to content and service creators.
1.2 The AiXpand vision
We are convinced that a true democratization of Artificial Intelligence would lower the costs of development and deployment of smart AI services while keeping the costs of processing at a minimum. Our vision is two-fold: (i) to convert computing devices to real assets and (ii) to democratize AI. We strongly believe that society will benefit immensely from transforming a wide range of computing devices, such as laptops, gaming stations, cloud infrastructure, consoles, and crypto-currency mining rigs, into real assets. These devices can then generate active income for their owners by executing jobs that have real societal and industrial value in a decentralized, peer-to-peer distributed fashion. Our AI democratization vision aims to bring AI use cases to end-consumers by reducing the go-to-market time, resources, and costs required by service providers, content creators, and software developers.
Our paper outlines an architectural approach to enable the democratization of AI, which aims to reduce the carbon footprints of complex systems and create ecosystems where AI application deployment relies on trustless sharing of processing power. The proposed approach also intends to reduce the time gap between the incubation of an idea and the launch of a marketable product. Our proposal builds upon existing research and development in the field of Machine Learning Operations, particularly upon the SOLIS[11] framework architecture. The additions include: decentralized job distribution layer, improved message security with internal blockchain consensus, and integration with EVM [12] compatible networks on-chain.
1.3 Summarizing our progress so far
In our previous work [11], we proposed an end-to-end architecture and methodology aimed at standardizing critical stages of production-grade machine learning pipelines, with a particular focus on edge-based systems. The primary goal was to offer technological independence, freedom to operate, and versatility in composing and deploying business rules. This resulted in an end-to-end MLOps framework that supports multiple jobs and workers, tensor framework agnosticism, and pre-built templates and serving pipelines for prominent frameworks such as TensorFlow [13] and PyTorch [14]. The currently proposed architecture enables seamless integration with a wide array of data sources, including CCTV cameras, relational databases, flat files, and sensor networks, while facilitating rapid deployment of business rules through low-code and no-code approaches. The component which lies at the foundation of this architecture, the AiXpand Execution Engine, can be easily configured for IoT-based protocols like MQTT [15] and AMQP [16], as well as web REST endpoints, with other endpoints implementable via a low-code plugin approach.
Building on the vision and objectives of our previous work, additional research and development have been carried out in the field of trustless job distribution within decentralized networks. Various peer-to-peer job distribution mechanisms and templates inspired by MapReduce[17] patterns have been devised. However, unlike traditional MapReduce approaches, a significant paradigm shift was necessary due to the trustless nature of the decentralized peer-to-peer network and the heterogeneity of the worker pool’s processing capacity and availability.
2 Related work
2.1 MLOps, DevOps and productization of AI
To the best of our knowledge, there has been no significant advancements in the field of Machine Learning Operations (MLOps) and Development Operations (DevOps) since our initial publication [11]. The focus of our research was to reduce the barriers to entry for the development of end-to-end AI-enabled systems or large-scale IoT-based applications. Classic DevOps methodologies are insufficient for the deployment of machine learning-based products due to the high computational demands. Therefore, it is necessary to optimize the use of available resources while maintaining a stable environment. Typically, machine learning applications require the use of both CPU and GPU, where the former is used for heuristics and business logic, while the latter is used for inference and data preprocessing/postprocessing. As a result, a specialized solution is required.
One such solution is Kubeflow [18], a framework-agnostic, open-source tool designed to streamline machine learning pipelines on Kubernetes. It supports most stages of the machine learning lifecycle, including data preprocessing but without data acquisition of any sort, model training and retraining, prediction serving and versioning, or service management. The main drawback of Kubeflow is that it requires extensive prior knowledge of Data Science and Kubernetes. Another drawback, that we are aware of, is that Kubeflow does not support full end-to-end application pipeline development but is rather focused only on the pure MLOps.
Another option is TensorFlow Serving [19], which allows for easy deployment of TensorFlow-based machine learning models through a low-code solution via a gRPC or HTTP endpoint. However, it does not provide tools for training, data acquisition, or business logic - not to mention highly complex end-to-end pipelines required in various types of applications.
ONNX [20] is an open format for representing machine learning models, regardless of the technologies used to build them. It currently supports conversions from popular machine learning frameworks such as TensorFlow[13], PyTorch[14], and SciKit-Learn, and its purpose is to provide a framework-agnostic interface for machine learning applications.
Finally, in relation to the AiXpand architecture, while the above-mentioned frameworks and technologies could be considered competing approaches for MLOps, most of them are in fact technologies that could be easily integrated and encapsulated within the AiXpand Execution Engine Pipelines.
2.2 From edge, fog & cloud to AI
Currently, there are three main approaches to providing AI-enabled applications: edge, fog, and cloud AI. Edge AI [21] involves deploying AI applications directly on the devices where data is generated. This approach allows users to utilize moderately advanced hardware solutions that are in close proximity to the data that requires processing. It is a cost-effective method that offers enhanced privacy. However, it lacks scalability when there is a growing demand for AI solutions. On the other hand, cloud AI enables vertical scaling by leveraging increasing computing power as needed. Nevertheless, it faces limitations at the network level when dealing with substantial amounts of data. Fog AI offers a horizontal scaling solution by utilizing a centralized network of edge devices. It aims to fulfill the high computing requirements of cloud AI while reducing overall network traffic. Both fog and cloud AI adopt subscription-based business models, which may not be advantageous for businesses that do not require real-time AI solutions. We propose a decentralized architecture that harnesses the computing capabilities of edge devices while maintaining low-latency connections between customers and AiXpand network nodes. Our architecture provides an on-demand, micro-transaction payment-based option for clients with AI-powered end-to-end application needs.
2.3 The Decentralized paradigm
2.3.1 The classic distributed computing
Distributed computing has turned into a cost-effective, high-performance and fault-tolerant reality due to great technological advances and falling costs of hardware. It refers to a collection of independent entities (workers), organized in a clear hierarchy with worker nodes and a central authority overseeing them. These workers computationally cooperate to solve a problem that would otherwise be potentially intractable at individual worker level. A distributed system is characterized, according to various authors [22], by the following features: (a) no common physical clock - which guarantees asynchronicity amongst the processors, (b) no shared memory - which means that message-passing is required for communication, (c) autonomy and heterogeneity - the processors are not part of a dedicated system but cooperate with one another to solve the problem jointly; they have different speeds and can be running a different operating system.
2.3.2 Transition to blockchain in distributed computing
Classic distributed computing systems are typically centralized and rely on a central authority to manage and process data. Whereas, Distributed Ledger Technology (DLT) [23] is a specific type of distributed computing that is characterized by its use of a decentralized and immutable ledger to record and share data across a network of nodes.
One of the key differences between distributed ledgers and classic distributed computing is the way data is stored and shared. In a distributed ledger, transactions are recorded on a public or private blockchain and then replicated across the network of nodes while in usual distributed computing scenarios data is centrally stored and processed. Thus, in decentralized setting. a tamper-proof record, which can be accessed by anyone on the network, of all transactions is possible while in the classic distributed case the centralized date is prone to unwanted modifications. Even more, in almost any implementation, distributed ledgers often use consensus mechanisms, such as those based on BFT[24], to ensure that all nodes on the network agree on the current state of the ledger. This guarantees that the data stored on the ledger is accurate and secure.
Therefore, while in distributed computing the system relies on a main or master processing entity in the decentralized setting a peer-to-peer organization replaces the server-workers architecture.
BitTorrent and IPFS (InterPlanetary File System)[25] are two examples of how traditional distributed computing systems are being integrated with blockchain technology to create decentralized, secure, and efficient solutions for data sharing and storage.
BitTorrent [26], one of the first decentralized protocols for file sharing, uses a peer-to-peer network to allow users to share files directly with each other, without the need for a central server. This greatly increases the speed and scalability of file sharing. In recent years, BitTorrent has been integrated with blockchain technology to create a decentralized file storage and sharing platform called BitTorrent File System (BTFS)[27]. BTFS allows users to share files and earn rewards in the form of cryptocurrency for contributing storage and bandwidth to the network.
Similarly, IPFS [25][28] is a peer-to-peer protocol for sharing files and data across the internet. It allows users to share files directly with each other, rather than rely on a central server. IPFS has also been integrated with blockchain technology to create a decentralized and distributed content delivery network called Filecoin [29]. Filecoin allows users to rent out their unused storage space and bandwidth to other users in exchange for Filecoin tokens.
2.3.3 AI & Blockchain
AI and blockchain technology can complement each other in a number of ways. One major benefit of using blockchain in conjunction with AI is the ability to securely and transparently share and access data. All actions, some of which could be expensive modeling, human data creation/annotation or even a simple AI-app-generated event, are traceable on a public distributed ledger. This can be particularly useful in areas such as medical research, where large amounts of sensitive data need to be shared among multiple organizations. Another benefit is that blockchain can be used to create decentralized platforms for AI model training and deployment, as described in a framework proposed by Microsoft Research[30]. This can help ensure that the models are fair and unbiased, as well as provide a way for individuals and organizations to share and monetize their own models.
A few examples of companies that are combining blockchain with AI include Ocean Protocol[31], that uses blockchain to create a decentralized marketplace for data, and SingularityNET[32], which fosters blockchain to create a decentralized platform for AI services.
Below, we list a few industries/sectors and benefits brought about by AI and blockchain.
Supply Chain Management: The combination of AI and blockchain can be used to create a more efficient and effective supply chain management system. For example, deep learning algorithms can be used to analyze data on supply chain transactions and optimize processes, while blockchain can be used to create a secure and transparent record of every transaction.
Financial Services: The combination of AI and blockchain can be used to create a more secure and efficient financial system. For example, deep learning algorithms can be used to detect and prevent fraudulent activities, while blockchain can be used to create a tamper-proof record of every transaction.
Healthcare: The combination of AI and blockchain can be used to create a more personalized and secure healthcare system. For example, deep learning algorithms can be used to analyze patient data and provide personalized treatment plans, while blockchain can be used to create a secure and privacy-preserving record of every patient’s medical history.
Energy Management: The combination of AI and blockchain can be used to create a more efficient and sustainable energy management system. For example, deep learning algorithms can be used to analyze data on energy production and consumption and optimize energy use, while blockchain can be used to create a transparent and auditable record of every energy transaction.
Decentralized Autonomous Organizations (DAOs): The combination of AI and blockchain can be used to create more effective and autonomous DAOs. For example, deep learning algorithms can be used to analyze data on user behavior and preferences and optimize product recommendations and investment decisions, while blockchain can be used to create a transparent and auditable record of every transaction within the organization.
2.3.4 Integrating AI on blockchain through smart contracts and microtransactions
Smart contracts and microtransactions are two key ways that blockchain technology can enable the integration of AI on the blockchain, by providing an infrastructure for automating transactions and for monetizing AI services. Smart contracts are self-executing contracts with the terms of the agreement written directly into code. They can be used to automate the execution of transactions when certain conditions are met, without the need for intermediaries.
This is particularly useful in the context of AI, as smart contracts can be triggered by AI systems to execute transactions automatically. For instance, an AI model may make predictions about the price of a stock. If the conditions set in the smart contract are met, the trade will be executed automatically.
Microtransactions, on the other hand, are small financial transactions that can be executed efficiently and inexpensively on the blockchain. They can be used to enable the monetization of AI services on the blockchain. For example, an AI model that is trained on a decentralized network could be made available to other users in exchange for small amounts of cryptocurrency. This would allow the creators of the model to earn income from their work, and it would also encourage the sharing and improvement of AI models.
3 Architecture
3.1 A top-down view
The proposed decentralized architecture, as illustrated in Figure1, is grounded in the well-established design of validator-like AiXpand Processing Nodes, which facilitate regional or global aggregation of AiXpand Processing Node Pools. Distinct from traditional validators responsible for blockchain protocol consensus, the primary function of an AiXpand Validator Node pertains to job validation and communication aggregation. Concurrently, all AiXpand Processing Nodes participate in a private blockchain consensus, ensuring secure and immutable job execution and data transfers.
Figure 1: Overall view of the AiXpand E2s network of networks
The overall AiXpand Ecosystem encompasses various middleware components, including IoT message broker tools and REST APIs for diverse use-cases. At its core is the AiXpand Execution Engine (E2), depicted in Figure 2, along with its platform-agnostic SDKs. The E2 can be deployed on a wide array of target operating systems and hardware devices, with or without advanced GPU compute capabilities. AiXpand employs both a private proprietary lite blockchain for internal use and an external public mainnet blockchain for microtransactions orchestrated by individual E2 processing nodes. Consequently, each E2 is uniquely identified and recognized within the network through a public blockchain (mainnet) non-fungible utility token. The subsequent paragraphs will discuss the primary decentralized processing distribution and the integration of a microtransaction-based worker reward system.
Although a comprehensive tokenomics strategy is proposed, it has not yet been publicly released. This paper focuses on the high-level technological aspects related to the AiXpand E2 Node Deeds’ interaction with the protocol, the fundamental principles of oracle-like communication between the mainnet blockchain (BC) and AiXpand E2s, and the overall algorithmic approach for reward distribution 321.
Figure 2: Top view of AiXpand Execution Engine ecosystem
In Figure1 the regional validation nodes coordinate the communication between nearby local nodes as well as provide inter-region distributed ledger data synchronization.
The Proof-of-AI protocol (PoAI) 1 manages the rewarding of users who operate AiXpand E2 processing nodes. This incentivizes the provision of compute power by both miners and service consumers. Miners use the AiXpand E2 solely for decentralized job execution, while service consumers use specific functionalities or applications developed by AiXpand content creators and service providers. As previously mentioned, an in-depth explanation of public on-chain algorithms and procedures is beyond the purpose of this paper, however the pseudo-code algorithmic approach of AiXpand Proof of AI (PoAI) can be analyzed in the current paper.
In order to summarize the funds allocation and distribution in our job execution agnostic reward pools we will denote: $T_{N}$ the total time a given node $N$, identified by a unique smart-contract based on a Node Deed in the overall population of all $A$ AiXpand Processing Nodes, is alive in the network; $T_{epoch}$ the protocol epoch time which is the interval used to calculate and distribute a round of rewards to all active nodes; $E$ the total number of epochs since the protocol genesys; $P_{N_{E}}$ the power score of a given node $N$ during epoch $E$. The final formula of the reward $R_{N_{E}}$ of the $Node$ $N$ at epoch $E$ is presented below in Equation5.
$T_{P}$ $=T_{epoch}*E$ (1)
$\tau_{N}$ $=\frac{T_{N}}{T_{P}}$ (2)
$I_{N_{E}}$ $=e^{P_{N_{E}}}*\tau_{N_{E}}$ (3)
$I_{E}$ $=\sum_{X\in A}{\frac{\tau_{X}*e^{P_{X_{E}}}}{T_{P}}}$ (4)
An important observation is that the power score $P_{i}$ for any node $i$ see equations 3 and 4 - can have both negative as well as positive values - i.e. the power score can get penalized in the negative range.
$R_{N_{E}}$ $=\frac{I_{N_{E}}}{I_{E}}$ (5)
The generic pseudo-code that depicts the proposed allocation of token rewards based on the previous equations 1,2,3,4 and 5 can be found in Algorithm 3. Worth mentioning is that, although ideally the proposed architecture relies on real online on-chain micro-transactions that incentivize the participants in a entirely trust-less fashion, the system can work without any problems in isolated environments. In such private, educational sandboxes or particularly isolated networks the reward tokens are treated as proofs of historical participation in the decentralized distributed processing.
3.2 End-to-end low-code pipelines
Developing end-to-end applications is a time-consuming process due to its multiple high-complexity stages. These stages include tasks such as developing and deploying communication modules and endpoints, implementing data acquisition streams, and developing advanced data processing techniques using AI-based algorithms or complex heuristics. Additionally, it involves establishing a continuous deployment strategy and executing it effectively. Usually multiple teams handle all these different software development streams using different skills and even different education backgrounds.
The AiXpand Ecosystem aims to simplify the process of developing and deploying complex application systems by dividing it into four distinct heterogeneous subsystems: communication, data acquisition, model serving, and business process as seen in Figure2. Each subsystem includes its own highly composable plugins, each with configurable parameters that allow customization of functionality without requiring additional code to be written. This versatility is achieved by allowing a pipeline to be configured by specifying which plugins to use and how they should be configured, enabling easy alteration of the pipeline’s functionality without code intervention or with an extremely low code footprint. This structure allows developers to use pre-existing or create custom low-code solutions for each of the subsystem within the network’s ecosystem, simplifying development, deployment and integration of complex application systems.
3.3 Secure execution with Plugin API
Due to the nature of trustless decentralized distributed job execution, multiple security issues arise: (a) custom code safety, (b) request and result message immutability, (c) data confidentiality and (d) solution proofing. Challenges (b), (c) and (d) are cryptographic issues that can be addressed either with the proposed distributed ledger architecture or with neural model training and inference based on homomorphic encryption. The former relies on code safety checking and signing.
Two primary types of business plugins have been developed, namely secure plugins and user plugins. These plugins are currently built using Python[33], a high-level general-purpose programming language. Business plugins can be considered either generic templates or default plugins for preparing business insights data/reports. On the other hand, secure plugins encompass a range of user plugins designed as proposals for future secure services, as well as code snippets created by anyone using the AiXpand E2 within the trustless network.
Both deployed user plugins and custom user code snippets are regarded as potentially unsafe due to their nature. Therefore, they are subject to limitations in terms of programming language keywords, functions, and packages. They must adhere to a specific higher-level Plugin API and undergo hash-sign verification, profiling, and code re-checking to prevent code injection and other tampering processes each time distributed execution is requested. This safety measure is implemented regardless of whether the code is transferred and executed either peer-to-peer or remotely by the same initiator and worker.
It is important to note that all these approaches can be accessed and utilized through external endpoints. This can be done either via a JavaScript/TypeScript-based SDK or through a Python client that utilizes our AiXpand E2 PyE2 package.
3.4 Multi-framework serving processes
The AiXpand Network and its ecosystem are designed to support applications powered AI. This is achieved by incorporating various AI models through the use of AI serving plugins within the serving subsystem. Additionally, the ecosystem is designed to accommodate applications that do not rely on AI models or techniques but are well-suited for decentralized environments.
To enable this, AiXpand provides a serving plugin interface, illustrated in Figure 3, which is not limited to specific frameworks like TensorFlow [13]. In this setup, all data input, inference, and prediction output are handled within the main thread of the main process. The implementation of the serving plugin interface primarily involves model loading, data pre-processing, and inference post-processing.
Consequently, the serving layer of the execution engine consists of multiple individual processes, with each process dedicated to running inference for a single model, whether it is a single-stage or multi-stage model. This allows for parallel execution of all the model-serving processes, ensuring that each process is encapsulated, easily monitored, and managed without interference.
Figure 3: High-level AiXpand E2 model serving processes UML diagram
3.5 Distributed decentralized execution
The architecture of the AiXpand Network is specifically designed to achieve scalable and decentralized distribution of different processing workloads. This is accomplished through a combination of peer-to-peer trustless job execution based on Map Reduce [17] and Scatter Gather [34] techniques, along with DLT. Throughout the design process, ensuring the security of all participants has been of paramount importance, influencing the network architecture and the distribution of processing workloads.
The ultimate objective is to empower a diverse range of participants, including service providers, work offload providers, and consumers, enabling them to fully leverage the capabilities offered by the network.
To provide a concise overview of the high-level process, we define an instance of the AiXpand E2 that requires external processing offload as a Main Node, while other nearby AiXpand E2 instances will be referred to as potential Worker Nodes. In this context, the Main Nodes have the capability to initiate distributed execution pipelines, each comprising two main processing stages, each with two distinct steps.
The initial stage, known as the Map Stage, is responsible for (i) searching for suitable nodes among all available Workers and (ii) mapping and assigning jobs to individual nodes. Following the Map Stage, the subsequent stage, referred to as the Reduce Stage, involves (iii) overseeing the completion of worker jobs and ultimately (iv) compiling and processing the results.
In order to achieve scalability within the trustless environment, it is crucial to address the security concerns associated with each stage (Map & Reduce) while maintaining coherence with the functional aspects of the distribution pipeline. To accomplish this, certain measures are implemented.
During the Map Stage, nodes are required to publish both their system specifications and job assignments to the blockchain. This ensures the immutability of the data. A Worker Node must sign and commit to its capabilities, as these factors impact its Power Score and, consequently, its share of profits.
In the Reduce Stage, Workers periodically send their job statuses along with Zero-Knowledge Proofs [35]. These proofs allow the Main Node to verify, with minimal computational effort, that overall progress is being made over time, without revealing sensitive information. This process ensures that the Main Node can assess the increasing progress of the tasks without compromising data privacy and security.
The primary goal of the AiXpand Ecosystem is to cater to a broad range of participants, including content creators, processing node operators, software developers, and end-consumers. To achieve this objective, AiXpand provides a comprehensive platform that is adaptable, reliable, secure, and easily scalable. This platform is designed to support not only AI-powered applications but also other decentralized applications, regardless of whether they utilize AI models or not.
Through its compatibility with various frameworks and the ability to execute multiple model-serving processes in parallel, AiXpand ensures an efficient and flexible solution for its users. The platform’s distinctive design, coupled with its integration with blockchain technology, enables secure and transparent operations, effectively meeting the diverse needs of participants.
To enhance the adoption of the ecosystem by the public, software development kits (SDKs) have been developed. These SDKs support multiple programming languages and frameworks, such as Python and NodeJS, significantly expanding integration options and improving the user experience. By using the SDKs, users gain access to the AiXpand Network through a partially functional paradigm, which offers callbacks for manipulating payloads received from AiXpand Nodes and provides simple yet powerful features for communication with the network. This empowers users to connect to the network, establish new end-to-end pipelines or join existing ones, create predefined or custom jobs within the pipeline, and receive notifications and heartbeats, facilitating seamless interaction with the ecosystem.
The AiXpand SDKs are intended to be used as a cohesive integration tool between the AI processing network and any existing or new software applications. They are designed to automate and manage the lifecycle of requests made on the AiXpand Network. AiXpand SDKs come in several flavors targeting either the Python or Node.js developers’ communities.
In the annex of this paper, we present a simple Python 3.x example 1 that demonstrates the process of connecting to a processing AiXpand E2 Node, executing tasks remotely, and distributing an end-to-end pipeline. However, it’s important to note that this example does not include the necessary steps to enable Proof-of-AI (PoAI) reward distribution, which involves compensating the processing node pool for the executed job.
3.5.1 The AiXpand SDKs
The AiXpand Node.js SDK consists of two main components. The first component aids in project setup and maintenance, handling tasks such as code generation and message validation using YAML schemas. The second component is responsible for runtime operations, serving as an intermediary for communication with the network and an example of it can be seen in example 2
To support developers in integrating network functionality into their solutions, a Command Line Interface (CLI) tool is provided. This tool assists in constructing the necessary schemas for communicating with deployed pipelines in the AiXpand Network. During project development or maintenance, developers can use the CLI to ensure that the command definitions for the pipelines are synchronized with the network.
The application runtime incorporates a package tailored for a specific programming language, which is meant to be seamlessly integrated into the business process services of the consuming application. The implementation details may vary depending on the programming language, but generally involve callbacks which trigger internal events. The responses obtained from these events are then analyzed and converted into data structures that are specific to the programming language in use. This allows developers to design and create business objects that align with the requirements and objectives of the application.
To incorporate AiXpand into a project, there are two approaches available. The first option is to execute a series of CLI commands using the AiXpand Node.js SDK. This will retrieve the necessary AI pipeline message schemas specifically designed for Node.js applications. Alternatively, for Python development teams, the AiXpand PyE2 SDK package can be installed, eliminating the need for CLI commands. Once this prerequisite is fulfilled, the imported libraries within the project will handle all validation and transformations between the application’s business objects and the expected message formats of the AiXpand Nodes.
To publish a message on the network, it is essential to establish a valid communication channel with a public AiXpand MQTT. This can be achieved by connecting to the AiXpand Validator Node or by using the local broker bundled with the AiXpand Node. Both options provide extensive availability of resources and published AI services. Furthermore, AiXpand PyE2 SDK provides built-in support for listening to payloads, intercepting errors, and diagnosing any issues that may arise due to incorrect use of the AI pipeline plugins.
3.6 Integration with Web3
The use of on-demand AI functionalities and capabilities offered by the AiXpand Network involves two main stages. The first stage involves the financial commitment of the processing power consumer, where a designated amount of tokens is allocated for covering AI request-related expenses. In the second stage, decentralized job execution takes place.
To facilitate efficient communication between local nodes, the AiXpand Network employs the AiXpand communication protocol, which utilizes MQTT as the underlying transport layer. MQTT is a widely adopted protocol known for its reliable scalability in IoT applications. The lifecycle of the AiXpand Protocol is encapsulated within the AiXpand PyE2 SDK, simplifying its implementation and usage.
To guarantee compensation for potential worker nodes’ efforts, it is essential that network requests are accompanied by a financial commitment. By including information pertaining to a valid inbound AiXpand wallet transaction, an AiXpand Node can process a designated workload. This step can be executed by the consumer application or via a separate procedure, as long as it is completed before submitting the workload to the network.
The AiXpand PyE2 SDK’s primary roles are to manage relevant financial information before disseminating the request on the network and to simplify the integration of this technology within external systems. For processes that necessitate multiple network requests, substantial transaction fees may be incurred. To minimize expenses related to utilizing cryptocurrency networks, depositing a larger sum into an escrow account is possible, allowing the AiXpand internal blockchain to oversee expended funds.
The AiXpand PyE2 SDK employs specific Web3 components for interacting with payment channels, contingent on implementation. On the AiXpand E2 side, a simple validation is performed to determine whether the request is signed by the AiXpand Node owner and run locally (in which case payment is unnecessary) or whether a valid funding transaction is attached.
The request entails disseminating a simple message to a user-controlled node or the nearest validator node. Subsequently, the network assumes responsibility for managing the request. Built with robust components, the AiXpand PyE2 SDK offers a reliable communication channel with the AI-enabled node network. Request progression is monitored by event listeners within the SDK, and any outcome is documented and relayed to the initiating processes.
Throughout pipeline execution, AiXpand Network Nodes generate status payloads, which outline the request’s current state, report potential errors, and supply information on generated assets. These payloads are monitored by event listeners within the AiXpand PyE2 SDK components, and appropriate handlers are provided for seamless interaction with ongoing processes.
3.7 AiXpand Rapid Application Development
The AiXpand Ecosystem envisions a future where AI-powered application development is accessible to individuals, regardless of their familiarity with programming languages. To achieve this, we are creating tools that enable both developers and non-developers to effortlessly articulate their business requirements and translate them into instructions compatible with the AiXpand SDKs.
For users with existing programming experience, our initial set of tools is designed to facilitate expressing their business needs. We provide code samples that demonstrate the functionality of the AiXpand Ecosystem. To streamline the process, we offer a code-sample gallery tool, allowing users to swiftly run samples, modify code and inputs, and test various results without needing to launch an IDE or create a new project.
To accommodate individuals without programming expertise, we are developing low-code tools that empower these users to generate their own code samples without extensive programming experience. Our ultimate goal is to provide a user-friendly interface that enables anyone, irrespective of their familiarity with source code and programming, to use drag-and-drop tools on-screen to express their desired business functionality.
AI-powered applications rely on data, and incorporating data from existing sources into applications is a crucial part of development. This typically requires previous programming experience. We aim at eliminating this requirement by providing visual connectors, allowing users to easily integrate existing data into their applications through a drag-and-drop graphical interface.
While multiple approaches already exist on the market, such as Azure ML [36] and Amazon SageMaker [37], these tools necessitate extensive knowledge of Data Engineering and Data Science. Our aim is to provide simpler visual means that will be usable in the decentralized ecosystem.
Regardless of one’s expertise, we are convinced that visual tools have the potential to greatly enhance the developer experience.
4 Applications
In this section of the paper, we present a range of applications that have been effectively implemented within the AiXpand Network. These applications cover diverse domains and industries, catering to both our internal usage and our customers’ requirements. Additionally, we have collaborated with external partners to develop applications that are specifically tailored for their needs.
While most of these use cases are presently deployed and operational across multiple private or public worker pools that utilize AiXpand E2 on GPU systems, the distribution of workload does not currently incorporate PoAI-based incentivization.
4.1 Physical security
The average human attention span is approximately 10 seconds[38], which decreases when an individual is tasked with monitoring multiple video streams simultaneously. In order to alleviate this strain, machine vision using deep learning models can be deployed in video surveillance to assist the operator in detecting unusual events. The multiple applications of this technology include industrial personal safety, location security, and video equipment security.
4.1.1 Industrial personal safety
This category of use-cases involves analyzing people’s posture and body positioning using a multi-stage machine learning model and potentially applying specific heuristics. By using a deep learning model to localize persons in an image and then a second model to determine the positioning of specific body parts, we can gather valuable data about the actors in the scene. Furthermore, using second stage deep classifiers missing or improper protection equipment wearing can be alerted. Finally, complex correlation can be made using multiple AI models - i.e. such as determining specific postures or dangerous situations in industrial areas where, for example, heavy machinery is operated.
Use-case: Fallen person
A problem that can occur because of the aforementioned attention issue, especially when the operator has to observe a bigger number of cameras, is the lack of immediate action taken when a person falls on the ground from various reasons and needs assistance. In some cases, a delay of even 5 minutes can make a substantial difference. As described above, we can use multi-stage models in order to ensure the safety of the working personal in a given environment. By analyzing the placing of some human joints and angles determined by those we can detect in a matter of seconds if an employee is lying on the ground and may need assistance.
4.1.2 Location security
This category involves analyzing the appearance of unusual elements or individuals in certain scenes, a use case specific for CCTV video analysis. To achieve this, we can use a deep vision detector that can be either class oriented - i.e. for detecting a animal in a place where no animals are supposed to be - or even class agnostic for detecting suspect objects.
Use-case: Vandalism
In this case a deep vision model is trained to detect either subtle or significant changes in a scene and thus can be used to identify signs of vandalism, such as graffiti on a wall, trash on the floor, moved furniture or even unknown objects dropped in the target scene. The integration of AI in video surveillance systems not only improves the efficiency of detecting vandalism, but also enables prompt intervention and mitigates potential security risks.
4.1.3 Video equipment security
While the previous category analyzes the scene in a given image, this third category analyzes the image itself in order to detect potential damage suffered by the recording equipment, either by natural causes (e.g. outdated cameras) or by human intervention (e.g. moved cameras). By analyzing the positioning of certain edges in the frame or other features provided we are able to quickly notify the technical staff in order to solve the issue at hand.
Use-case: Camera quality check
In this case a deep vision model is employed to identify minor deviations in the placement of edges or abrupt fluctuations in frame entropy, thereby detecting any potential damage or malfunction in the video cameras. By leveraging these features, the model is able to promptly alert technical personnel to any issues that require immediate attention. This use-case not only guarantees the dependability and efficiency of the surveillance system, but also saves valuable time and resources that would otherwise be spent on manual monitoring and equipment inspection.
4.2 Predictive analytics
Predictive analytics is the use of data, statistical algorithms and machine learning techniques to identify the likelihood of future outcomes based on historical data [39]. Such techniques are used by a large variety of businesses - from retail brick and mortar stores to physical security companies in order to optimize their business processes or to make predictions regarding the future behaviour of people or equipment.
4.2.1 Replenishment
One of the biggest problems of retail logistics is that of over-supply and under-supply. Over-supply can incur losses based on the perishability of the over-supplied products, as the stock is not sold in time or based on the fact that storage is also a cost; while under-supply may induce missed potential sales due to the product not being in stock.
Use-case: Predictive replenishment
The easiest, but not the most efficient solution, is the use of heuristics [40] to generate replenishment orders. The main drawback is that good results require very complex algorithms which are hard to maintain and do not easily scale. The alternative to the heuristic method is the automated learning approach. Machine learning techniques, especially deep representation models [41], are able to more accurately find patterns in historical data and using continuous training methods can easily adapt to new circumstances in order to provide reliable consumption predictions.
4.2.2 Stock prediction
Traditionally, the user interaction with a product is an integral part of the product itself. Therefore, significant effort is dedicated to ensuring a streamlined user journey where all available choices are visible and easily discernible. This involves employing established user experience patterns for menus, buttons, layouts, and other components that facilitate the manipulation of data within a product. AI presents a distinctive challenge. The AI product itself cannot be directly experienced, as users can only interact with the resulting output — an amalgamation of vast amounts of data. This characteristic allows for seamless integration of AI into other products.
Use-case: Stock prediction in a spreadsheet
An intriguing application of this fusion between AI and everyday operations is the inclusion of prediction functionality within a spreadsheet. To illustrate this application, we can demonstrate how the capabilities of the AiXpand Network can be leveraged. A user selects a specific range of cells within the spreadsheet that contains relevant values for the desired prediction. These values are then transmitted to an API that incorporates AiXpand Nodes. The API performs validation on the received data, formats it for communication with the nearest AiXpand Node, and awaits the response generated by the predictive model. Finally, the response is relayed back to the spreadsheet for display.
The communication between the spreadsheet and the API can be achieved through a RESTful architecture or in real-time using a duplex channel such as websockets. The API utilizes the AiXpand Node.js SDK and taps into the predictive models published within the network to generate the results that are then relayed back to the user of the spreadsheet. For a more detailed depiction of the underlying processes, refer to the swimlane diagram provided in Figure4.
Figure 4: AiXpand network integration in a Google Sheets plugin
4.2.3 Use-case: Predictive Maintenance
A prominent use-case for deploying predictive analytics applications within decentralized infrastructures is industrial predictive maintenance. Predictive maintenance refers to the proactive identification and resolution of potential failure points in industrial equipment before critical malfunctions transpire [42]. This is typically achieved by monitoring equipment behavior changes and comparing them to the historical behavior of similar items.
Traditional solutions for predicting equipment breakdown, such as autoregressive models like ARIMA [43], often suffer from low accuracy and an inability to detect subtle changes in equipment behavior. Consequently, deep learning-based methods [44] are employed to create accurate and scalable systems. These deep models can effectively recognize patterns, predict future malfunctions, and readily generalize to new situations, offering a reliable solution.
4.2.4 Use-case: Anomaly Detection
Closely related to predictive maintenance, anomaly detection is another crucial use-case. Anomaly detection involves the identification of data points that significantly deviate from the majority of the data [45]. It can be applied in physical security applications to detect potential security threats or system malfunctions. Such systems are essential due to the high volume of telemetry and events, which are costly to parse by human operators and prone to errors due to the tedious nature of the task. A common solution to the anomaly detection problem is the use of Multivariate Gaussian Models to identify suspicious events [46]. While this method has low computational cost and is easy to implement, it suffers from low accuracy and an inability to represent and leverage complex data patterns. Alternatively, deep representation models [47] can be employed to take advantage of higher-order patterns, enhancing the accuracy and effectiveness of anomaly detection systems.
4.3 Video Anonymization - Motionmask™
Motionmask™ illustrates the integration of the AiXpand Network within the backend infrastructure of a traditional Web2 service. The application serves as a web-based tool for anonymizing videos and ensuring compliance with data protection regulations through the provision of an user-friendly interface for identifying and masking specific individuals or vehicles. In order to have a clear view, the swimlane diagram from Figure5 describes the underlying processes in more detail.
Figure 5: AiXpand Network integration in Motionmask™information flows
At the core of Motionmask™ lie very powerful AI detection models. These models require a lot of computing resources which, since Motionmask™ is an online available tool, would have to be provided solely by Motionmask™ itself. This would not be possible and efficient without the AiXpand Network because it would require for Motionmask™ to have an infrastructure small enough to avoid wastage, but also big enough to reliably support surges in user requests.
The interface, which operates within a web browser, communicates with the backend by making RESTful API calls. The backend, in turn, manages all the application business logic such as handling user creation and deletion, registering requests for video anonymization and handling file uploads. The integration of the AiXpand Network is done within a dedicated internal service that connects to the provided MQTT broker and communicates with the nearest AiXpand Node.
Progress is tracked by witnessing the events emitted by the AiXpand Node which describe the state of the running pipeline. Upon completion of processing, the backend services publish an event that is propagated back to the user interface, and the resultant output is presented to the user.
The payment for services provided by the Motionmask™ application can be completed using both traditional payment processing services and cryptocurrency. Specifically, payments can be made through traditional payment processing channels with FIAT currency or through the transmission of tokens to the Motionmask™ smart contract.
4.4 Other applications
4.4.1 Decentralized game engine
In the previous examples, the primary focus was on use-cases that employ artificial intelligence particularly shallow and deep machine learning models. Nevertheless, the ecosystem is also capable of supporting the development, integration and even migration of non-AI applications given the hypothesis that the particular target application is suitable for running in a decentralized environment. To demonstrate this versatility, a simple sandbox simulation life-game was designed showcasing how virtual agents interact with one another in an evolutionary heuristic environment. While the initial purpose of this sandbox game was to illustrate the flexibility of AiXpand Development Ecosystem, this mini-project can be used as a tutorial or even as a basic application template.
While being based on the architecture presented in Figure 2, most of the development effort has been geared towards three areas: (i) sandbox simulation encapsulated as a AiXpand E2 model serving module,(ii) the business logic that prepares the pre-rendering vector data and finally (iii) the UI that renders and enables interaction with the sandbox game. The model serving component implements the life-game simulation stateful process itself instead of a traditional stateless machine learning model. While in a conventional model serving process the inferred data is based on input data generated from an upstream data acquisitor, the simulation process here produces the state of the environment after each cycle. The business heuristics component is then responsible for preparing the data that will be delivered from AiXpand E2 to an external custom user interface. Furthermore, the custom UI consumes the inflow of data while generating interaction events in turn that are delivered to the AiXpand Processing Node which runs the simulation game.
5 Ongoing work
Our development teams are currently working on several crucial areas aimed at addressing key aspects of decentralized execution. These areas include optimizing AI inference and training processes by leveraging homomorphic encryption through our AiXpand EDIL Framework (EDIL - Encrypted Decentralized Inference and Learning) . Our framework currently supports popular deep learning frameworks such as PyTorch and TensorFlow. Additionally, we are actively exploring the utilization of zero-knowledge proofs for evaluating job completion and refining essential workflows involving two-blockchains. This includes establishing efficient interactions between on-chain systems and the AiXpand E2 environment through oracles-based mechanisms. These ongoing developments demonstrate our commitment to advancing decentralized execution capabilities and enhancing the overall functionality and security of the AiXpand Ecosystem.
5.1 Distributed Homomorphic Encryption
In decentralized MLOps, ensuring secure and confidential distribution of jobs and data manipulation is essential for maintaining privacy. Traditional encryption techniques that rely on cipher keys require the disclosure of original data to authorized parties, which is not ideal in a distributed system aiming to provide complete anonymization. To overcome this challenge, our proposed framework, known as EDIL, introduces distributed homomorphic encryption (DHE). By leveraging DHE, we enable encrypted data processing without the need for decryption, ensuring both data confidentiality and security.
One of the significant advantages of DHE is its ability to allow multiple parties to perform computations on the same encrypted data without revealing any information to each other. This characteristic makes it suitable for a trust-less environment where sensitive data must be managed securely without any disclosure.
Considering the computational demands of MLOps jobs, it is crucial to minimize the overhead introduced by homomorphic encryption to ensure scalability. To address this concern, our proposed DHE approach utilizes a Domain Auto-encoder-based method, which reduces the amount of information needed while retaining essential data features. From the perspective of a AiXpand Worker, the encoder functions as a one-way function without the decoder, preserving the confidentiality of the data. However, to maintain minimal overhead and achieve the desired secrecy, a unique auto-encoder is required for each job type. This approach allows for secure and efficient manipulation of sensitive data within a decentralized ML environment.
5.1.1 Decentralized Federated Learning
To dispatch a training job to AiXpand Worker Nodes, the self-designated AiXpand Main Initiator Node either automatically trains a DHE domain auto-encoder with the data to be processed or uses a predefined one from an existing library within the AiXpand Ecosystem. This generates an encoder compressing the data’s salient features. The Initiator Node can then define the end-to-end training job, including the machine deep learning model that consumes the domain encoder output and sends the compressed and DHE data to each worker-peer. Subsequently, workers utilize this DHE approach to train the model using the encoded information in multiple rounds, similar to FedAvg [48], under the supervision of the AiXpand Initiator Node.
5.1.2 Encrypted Inference
For deploying an inference job to AiXpand Nodes, the AiXpand Initiator Node or external API first encodes the raw data using a neural model stub or a pre-trained neural encoder in compliance with the proposed DHE approach. The Initiator Node then sends both the Main neural model and the encoded data to the partner nodes accepting the job. The receiving nodes forward the input data through the distributed Main model and execute specific post-processing heuristics and code required by the Initiator Node before sending the final result. Throughout this process, receiver nodes only have access to the encoded format received from the Initiator Node as proposed by DHE. By employing this approach, computational load can be distributed across a peer-to-peer decentralized ecosystem.
5.2 Two layers of BC
The AiXpand Network utilizes a dual blockchain system, consisting of a public and private blockchain, to derive optimal benefits from micro-transactions in the financial domain as well as implement distributed data transactions immutability, respectively. It is imperative to note that while a single blockchain can potentially achieve these objectives, this approach is not favored as it would result in an undesirable gas cost associated with the commitment protocol for data-only transactions that should accommodate an arbitrary load with no correlated financial burden.
The private blockchain is designed as a proprietary, lightweight system to guarantee operational freedom and minimize its architectural footprint. This blockchain can be employed in isolated, private settings and serves the primary function of enabling synchronization among the participating nodes in the network. This synchronization is crucial for the execution of delegated tasks and the signing of system specifications, both of which are integral to the operation of the AiXpand distributed jobs. These processes enable trust-less operations in a heterogeneous operating environment as we aim to achieve a fully trust-less ecosystem for job execution and rewarding.
In contrast, the public blockchain relies on the existence of the private blockchain and is unable to operate independently. The primary objective of the public blockchain is to provide a straightforward yet powerful interface to facilitate the monetization of the entire system through a decentralized approach. This is achieved through the integration of an oracle between the public and private blockchain, which connects payments to job requests and the progress of those jobs. To further safeguard the integrity of the system, the public blockchain incorporates a democratic mechanism to deter fraudulent activities and promote transparency. For instance, participants within the public blockchain can issue challenges to one another in regards to job quality, with the outcome of these challenges resolved through a random selection of nodes, and recorded immutably on the blockchain for future reference. The challenge mechanism incurs costs supported by the challenger until their accusations are proven true, at which point they recover their funds.
5.3 Oracles between on-chain and AiXpand E2 worlds
To attain the required level of automation within the ecosystem for multiple aspects from micro-transactions to node reliability history, the deployment of multiple oracles is necessary for the facilitation of communication between the on-chain and AiXpand E2. The oracles enable off-chain to on-chain data transfer between the private and public blockchains in order to enable the allocation of jobs toAixpand Worker Nodes and the creation of worker incentivisation escrow pools. Furthermore, they enable sending commands to worker nodes and retrieve results, assist in computing proofs of job execution and enable on-chain publishing of the outcomes. As central components of the automated end-to-end process, oracles must be secured to prevent the occurrence of vulnerabilities that may have severe implications for the system. It is therefore of utmost importance to ensure that oracles behave in a fair and correct manner, even in scenarios where attackers attempt to compromise the system.
5.4 Zero Knowledge Proof
In order to enable trust-less data consumption and jobs within the decentralized distributed job execution ecosystem, it is crucial to implement a minimal Zero Knowledge Proof [35] framework and capability in the system. The use of zero-knowledge proof allows the verification of the progress made by the AiXpand Worker Nodes without disclosing any sensitive information regarding the progress of the job, where one only needs to verify that it is increasing, as well as allowing the job initiator to verify the actual completeness of a given finalized job in a feasible time. This not only maintains the confidentiality of the processed data but also minimizes network traffic by reducing the payload exchanged between the Worker and Main Nodes. In this way, the proposed system aims to provide users with a secure and efficient network for the execution of various types of distributed jobs.
5.5 Conclusion
The vision of the AiXpand revolves around democratizing AI-powered application development and deployment as a transformative ecosystem to address multifaceted challenges and facilitate innovative solutions. All of our tools and technologies streamline the process of converting business requirements into instructions, enabling individuals with diverse backgrounds to seamlessly undergo this transformation.
In a broader context, the AiXpand Ecosystem embodies the remarkable potential of blockchain technology in establishing a more efficient, and inclusive ecosystem, extending beyond the realm of AI. Through its decentralized nature, the ecosystem strives to promote greater autonomy, resilience, and equitable participation, offering a promising paradigm for future advancements in various domains.
AiXpand serves as a testament to the potential of converting computing devices to real assets, while being a pioneer of AI democratization.